# All Optical Echo State Network Reservoir Computing

### v1

@@
All Optical Echo State Network Reservoir Computing
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
We propose an innovative design for an all-optical Echo State Network (ESN), an advanced type of reservoir computer known for its universal computational capabilities. Our design enables fully optical implementation of arbitrary ESNs, featuring complete flexibility in optical matrix multiplication and nonlinear activation. Leveraging the nonlinear characteristics of stimulated Brillouin scattering (SBS), the architecture efficiently realizes measurement-free operations crucial for reservoir computing. The approach significantly reduces computational overhead and energy consumption compared to traditional software-based methods. Comprehensive simulations validate the system’s memory capacity, nonlinear processing strength, and polynomial algebra capabilities, showcasing performance comparable to software ESNs across key benchmark tasks. Our design establishes a feasible, scalable, and universally applicable framework for optical reservoir computing, suitable for diverse machine learning applications.
1 Introduction
@@
Optical platforms for computation and machine learning can, in principle, achieve more efficient computation in terms of energy and time [13, 14]. One important challenge in all-optical approaches is how to realize the nonlinear activation. For this, various nonlinear optical processes have been studied including the nonlinear behavior of SBS [15]. It is important to note that, while the implementations of all-optical reservoir computers have been reported [16], the realization of an all-optical RC with universal computation capability (i.e., universality) has not yet been demonstrated.
@@
Here, we propose an innovative way to realize an all-optical implementation of any desired ESN configuration. The advantages provided by an all-optical architecture, as stated earlier, are fast computation and better energy efficiency than large software alternatives. For this reason, there has been growing interest in recent years in optical implementations of the matrix multiplication stage of machine learning architectures [19, 20, 21]. Here, we propose an in-fiber method for optical matrix-vector multiplication, which is followed by an in-fiber solution for optical nonlinear activation inspired by Slinkov et. al [22] using SBS. This unique combination of measurement-free matrix multiplication and nonlinear activation allows for a feasible design of the optical ESN.
The paper is organized as follows: Section 2 covers the basic concepts of ESNs and the theory behind the SBS linear amplifiers and activation function, Section 3 addresses our proposed design for the all-optical ESN, and Section 4 compiles the performance of the ESN on various tasks that evaluate its universality, and compares it to a software ESN of the same size. The tunability of the nonlinearity of SBS interaction is also analyzed in this section. Finally, Section 5 contains the concluding remarks.
2 Optical Reservoir Computing Concepts
2.1 Echo State Network
@@
2.2 Nonlinear Activation Mechanism
@@
The behavior of the transcendental Equation (7) has been simulated numerically, and is shown in Figure 1. Particularly, the nonlinear behavior of SBS looks very similar to a sigmoid function, which is often used as an activation function for machine learning applications. The linear behavior is tunable by varying the pump power and the nonlinear regime shows different behavior for different values of the $gL$ product given in the exponent of Equation (7), and also for different values of the pump splitting ratio. This allows for an optimizable activation function, which we expand on later in the paper.
(a) SBS Amplifier
(b) SBS Activation Function
Figure 1: (a) SBS operating in large pump (linear) regime. Input-output relationship is shown for various pump powers to demonstrate tunability. (b) SBS operating in self-pumped (nonlinear) regime. The nonlinear shape can be designed by choosing an appropriate $gL$ value.
3 Optical ESN Model
@@
The architecture aims to physically realize the ESN update Equation (1) at each time step by sequentially performing matrix-vector multiplications followed by nonlinear activation Equation (4). As shown in Figure 2, the input $u_{k}$ is encoded by a variable attenuator (VOA) at each time step $k$. The required $B^{\prime}$ vector elements are encoded using Brillouin amplifiers with appropriately tuned pump powers according to Equation (7) operating in the large pump regime, which gives a linear input-output relation as shown in Figure 1. The vector $B^{\prime}$ is chosen such that $AB^{\prime}=B$ to satisfy the ESN design in Equation (4).
The input vector is combined with the optical reservoir using fiber beam combiners. The matrix multiplication is performed in a series of sub-steps that each time step $k$ is divided into. Each sub-step is denoted by the index $m$, and represents multiplication of the elements of the $x_{k}$ vector with each successive row of the matrix $A$. As detailed above, these multiplications are easily performed by appropriately tuning the pump powers in the array of Brillouin amplifiers. The output from this matrix multiplication on the current sub-step is passed through a compensatory Brillouin amplifier to account for any losses during previous and successive fiber splitting and combining steps, and to pre-amplify the signal before the activation step. It is then passed to the nonlinear activation.
@@
The input to the activation function Brillouin amplifier must be used in a self-pumping scheme [22]. The previous amplification step allows for it to have sufficient power to pump the nonlinear interaction, giving a nonlinear function at the output as shown in Figure 1. Here, we also choose the ratio between the pump and input beams (the pump splitting ratio) appropriately to optimize the nonlinear shape of the activation function, which is expanded on in the next section.
The output is the $m$’th element of the state of the reservoir at the next time step, $k+1$. It is switched into the appropriate output branch by a $1:N$ tunable switcher and the output is given an appropriate delay and split before both being sent to a detector and back into the reservoir for the next time step. This design utilizes only a single nonlinear activation element, allowing for high scalability.
@@
After the necessary delays to line up all of the output vectors, and at the end of all of the sub-steps, the state vector has been successfully updated and fed back into the reservoir for the next time step. After collecting all of the state vectors in the detection steps, we use post-training according to Equation (2) to achieve the required estimated output in response to the time-series inputs $u_{k}$.
@@
This design allows for the choice of arbitrary matrix elements for the matrix $A$ and vector $B$. If chosen according to the convergence properties of an ESN, this design can be used to implement an arbitrary ESN, which is a universal family of reservoir computers [17]. This means that given a desired accuracy for a task (such as prediction), this method can be used to realize an optical hardware that achieves that accuracy. The proposed design also allows for scalability, as the number of nodes is increased simply by increasing the number of parallel fibers running in the optical fiber reservoir. Although, scalability is naturally limited by available resources and energy.
However, there is a trade-off between scalability and speed of operation involved. The current approach uses time multiplexing to simplify the hardware requirements and allows us to implement the nonlinear activation with only a single module. This means that with an increasing number of nodes, which may be required to achieve arbitrarily high accuracy for a given task, the number of time-steps required per successful loop will increase, which decreases the speed of operation.
Without this compromise, the hardware requirements grow in complexity and become difficult to scale, requiring $N^{2}$ linear Brillouin amplifiers and $N$ nonlinear modules. Thus, we have decided to allow for a reduced speed of operation to ensure a scalable and relatively simpler hardware design.
4 Results and Discussion
@@
Good performance in machine learning hinges on the ability to approximate nonlinear input–output relationships and capture complex temporal dynamics. According to the Stone–Weierstrass theorem [26], the polynomial algebra property is central to achieving universality. Therefore, we selected three distinct tasks to evaluate these critical capabilities. They are the sine-square classification test, prediction of the NARMA10 system, and the Mackey-Glass time series prediction. Each task evaluates mainly the reservoir computer nonlinearity, polynomial algebra capabilities, and dynamic memory retention, respectively. Additionally, we address the nonlinear properties of the SBS interaction and, thus, the optimization of the activation function. To achieve good results in the tasks mentioned above, we require strong enough nonlinearity, while maintaining the convergence conditions of the ESN. We vary the value of the $gL$ product as well as the pump splitting ratio with these constraints in mind to optimize the performance of the ESN in the selected benchmark tests. These results and investigations are further expanded on below.
@@
(a) Software (Sigmoid) Sine Square Classification, NMSE=0.14.
(b) Optical Sine Square Classification,
NMSE=0.12.
For each test, we analyze the effect of the nonlinear strength of the SBS interaction, given by the $gL$ product value of the medium and the splitting ratio for the self-pumping, which refers to the ratio of the pump power to the input power. Here, $g$ represents the Brillouin gain and $L$ represents the interaction length. These values appear in Equation (7). The $gL$ product decides the nonlinear shape and the parking point of the activation function, as shown in Figure 1. The pump splitting ratio independently changes the parking point of the activation function, which allows us to optimize it for greater performance. We also evaluate the performance as the number of nodes $N$ is increased, which shows the effect of scaling the ESN. These behaviors are task-dependent and the exact values of these parameters are subject to change with randomization of the ESN and inputs but, by extracting a range of top-performing parameters across the board, we attempt to design the most generally applicable SBS activation function and optical ESN. These investigations are expanded below.
4.1 Sine Square Classification
@@
(a) Average NMSE vs. Pump Splitting ratio for various $gL$, $N=10$.
(b) Average NMSE vs. Pump Splitting ratio for various $N$, $gL=12$.
(c) Average NMSE vs. $gL$ values for various $N$, $\mathrm{Pump/Input}=1200$.
@@
We evaluate the effect of varying activation function parameters on the NMSE for this test. We have computed the average NMSE over 250 variations of the $A$ and $B$ matrices to see clear trends in the NMSE. From Figures 4 (a) and (b), NMSE decreases consistently with increasing pump splitting ratios, indicating improved nonlinear activation efficiency. Figure 4 (c) shows that for a fixed pump splitting ratio as well, there is an optimal value for the $gL$ product in the range of $3-7\;W^{-1}$. We see that for $N=4$ nodes, the trends are difficult to see and the NMSE values are much higher. This is because the fitting is difficult for a smaller number of nodes. Therefore, we find for this task that $N=5$ and above are required for reasonable performance, with an increase in $N$ taking the performance closer and closer to saturation near $\mathrm{NMSE}\approx 0$. We also see, from Figure 4 (a) that there is a different optimum point in the performance for each $gL$ value. The most probable reasoning for this behavior is that both the pump splitting ratio and the $gL$ value are together responsible for the optimal parking point of the activation function, resulting in local minima.
The values reported by Slinkov et. al [22] allow the possibility of $gL$ values from $3-7$ with fiber lengths of $3-6$m at reasonable threshold powers of $1.26$ W which are very practical lengths of fiber. Experimental demonstrations have shown much higher levels of nonlinearity in the right materials [28, 29].
@@
4.2 NARMA10

### v2

@@
Abstract
We propose an innovative design for an all-optical Echo State Network (ESN), an advanced type of reservoir computer known for its universal computational capabilities. Our design enables fully optical implementation of arbitrary ESNs, featuring flexibility in optical matrix multiplication and nonlinear activation. Leveraging the nonlinear characteristics of stimulated Brillouin scattering (SBS), the architecture efficiently realizes measurement-free nonlinear activation. The approach significantly reduces computational overhead and energy consumption compared to traditional software-based methods. Comprehensive simulations validate the system’s memory capacity, nonlinear processing strength, and polynomial algebra capabilities, showcasing performance comparable to software ESNs across key benchmark tasks. Our design establishes a feasible, scalable, and universally applicable framework for optical reservoir computing, suitable for diverse machine learning applications.
I Introduction
@@
Optical platforms for computation and machine learning can, in principle, achieve more efficient computation in terms of energy and time [13, 14]. One important challenge in all-optical approaches is how to realize the nonlinear activation. For this, various nonlinear optical processes have been studied including the nonlinear behavior of SBS [15]. It is important to note that, while the implementations of all-optical reservoir computers have been reported [16], the realization of an all-optical RC with universal computation capability (i.e., universality) has not yet been demonstrated to our best knowledge.
@@
Here, we propose an innovative way to realize an all-optical implementation of any desired ESN configuration. The advantages provided by an all-optical reservoir architecture, as stated earlier, are fast computation and better energy efficiency than large software alternatives. For this reason, there has been growing interest in recent years in optical implementations of the matrix multiplication stage of machine learning architectures [19, 20, 21]. Here, we propose an in-fiber method for optical matrix-vector multiplication, which is followed by an in-fiber solution for optical nonlinear activation inspired by Slinkov et. al [22] using SBS. This unique combination of measurement-free matrix multiplication and nonlinear activation allows for a feasible design of the optical ESN.
The paper is organized as follows: Section II covers the basic concepts of ESNs and the theory behind the SBS linear amplifiers and activation function, Section III addresses our proposed design for the all-optical ESN, and Section IV compiles the performance of the ESN on various tasks that evaluate its universality, and compares it to a software ESN of the same size. The tunability of the nonlinearity of SBS interaction is also analyzed in this section. Finally, Section V contains the concluding remarks.
II Optical Reservoir Computing Concepts
II.1 Echo State Network
@@
(a) SBS Amplifier
(b) SBS Activation Function
(c) Sigmoid Function
Figure 1: (a) SBS operating in large pump (linear) regime. Input-output relationship is shown for various pump powers to demonstrate tunability. (b) SBS operating in self-pumped (nonlinear) regime. The nonlinear shape can be designed by choosing an appropriate gLgL value. (c) Sigmoid function for comparison.
II.2 Nonlinear Activation Mechanism
@@
The behavior of the transcendental Equation (7) has been simulated numerically, and is shown in Figure 1. Particularly, the nonlinear behavior of SBS looks very similar to a sigmoid function, which is often used as an activation function for machine learning applications. The sigmoid function is plotted for comparison in Figure 1(c). The linear behavior is tunable by varying the pump power and the nonlinear regime shows different behavior for different values of the $gL$ product given in the exponent of Equation (7), and also for different values of the pump splitting ratio. This allows for an optimizable activation function, which we expand on later in the paper.
III Optical ESN Model
@@
The architecture aims to physically realize the ESN update Equation (1) at each time step by sequentially performing matrix-vector multiplications followed by nonlinear activation Equation (4). As shown in Figure 2, the input $u_{k}$ is encoded by a variable attenuator (VOA) at each time step $k$. The required $B^{\prime}$ vector elements are encoded using Brillouin amplifiers with appropriately tuned pump powers according to Equation (7) operating in the large pump regime, which gives a linear input-output relation as shown in Figure 1. The vector $B^{\prime}$ is chosen such that $AB^{\prime}=B$ to satisfy the ESN design in Equation (4). The values of the $A$ matrix are realized with a combination of a VOA and a Brillouin amplifier. This allows access to arbitrary (positive or negative) matrix elements, as discussed further below.
The input vector is combined with the optical reservoir using fiber beam combiners. These are coherent optical beam combiners, which introduces the possibility of reduced performance due to phase fluctuations. Fiber systems are known to have phase fluctuations due to vibrations, which necessitates the use of a phase lock loop to ensure coherent operation of the beam combiners. Electronic servo systems, such as the system proposed by Liang et al. [26] can be used to mitigate this issue.
The matrix multiplication is performed in a series of sub-steps that each time step $k$ is divided into. Each sub-step is denoted by the index $m$, and represents multiplication of the elements of the $x_{k}$ vector with each successive row of the matrix $A$. These multiplications are performed by appropriately tuning the array of VOAs for the inputs and the pump powers in the array of Brillouin amplifiers. As these amplifiers must be modulated every sub-step, the pump powers are also controlled by VOAs. We allow for negative elements in the ESN design matrices by representing the values of the state vector in dBm units.
While the Brillouin amplifiers implement only positive multiplication values (as the operation takes place in units of W), the accompanying VOAs can be tuned in each sub-step to achieve the desired negative value at the output in dBm units. This allows us to access negative matrix elements by using the appropriate range of optical powers below 0 dBm. Thus, we realized the ESN equations (1) - (4) all in dBm scale. This method is possible as the output is continuously monitored by the final detection step and the Brillouin amplifiers must be updated multiple times per input time step in order to time-multiplex the values of the $A$ matrix. While the elements of the $A$ and $B$ matrices are fixed constants, the amplifier gains and VOA losses are adjusted in each instance to convert the equations from dBm units into physically realizable gain and loss values.
As an example, if we have a specific input to output relation for this matrix multiplication given by $s_{i}=A_{ij}x_{j}$, where $x_{j}$ is 1-1 dBm and we wish to multiply $A_{ij}=2$, we would set our Brillouin gain to no gain, and set the VOA to attenuate the signal in order to achieve an output $s_{i}$ of 2-2 dBm. If the input $x_{i}$ is +1+1 dBm and we wish to multiply still the same value $A_{ij}=2$, we would set the VOA attenuation to no attenuation, and set the Brillouin gain to amplify the signal to make $s_{i}=+2$ dBm. The gain and attenuation values are calculated as Gain or Loss $10\mathrm{log}_{10}(\frac{P_{\mathrm{out}}}{P_{\mathrm{in}}})$. We can control the Brillouin gain by appropriately tuning the pump power, as explained above. In this way we are able to time multiplex the construction of the $A$ matrix by continuously monitoring the input and varying the VOA attenuation and Brillouin gain accordingly.
At first glance, this might seem like a limitation. However, the conversion from matrix elements in dBm scale to actual hardware parameters is straightforward and does not significantly increase implementation difficulty. We have accounted for this in simulations as well, as is discussed further in Section IV. This also requires the SBS process to respond quickly enough to modulations of the pump. The realistic gain modulation speed of fiber Brillouin amplifiers by modulating the pump power is typically limited to tens of MHz (e.g., up to around 50–100 MHz in silica fibers). This limitation arises from the intrinsic response time of the SBS process, which is governed by the phonon lifetime (approximately 10 ns) or equivalently the Brillouin gain linewidth (50–100 MHz). For example, recent advances in optical imaging [27] have made use of the quick response time of Brillouin amplifiers by modulating at over 1 MHz. Modulating the pump amplitude faster than the limit set by the Brillouin gain linewidth would not allow the acoustic wave (phonons) to fully respond, resulting in reduced or ineffective gain modulation. This tradeoff is discussed further below.
@@
The output from this matrix multiplication on the current sub-step is passed through a compensatory Brillouin amplifier to account for any losses during previous and successive fiber splitting and combining steps, and to pre-amplify the signal before the activation step. It is then passed to the nonlinear activation. The input to the activation function Brillouin amplifier must be used in a self-pumping scheme [22]. The previous amplification step allows for it to have sufficient power to pump the nonlinear interaction, giving a nonlinear function at the output as shown in Figure 1. Here, we also choose the ratio between the pump and input beams (the pump splitting ratio) appropriately to optimize the nonlinear shape of the activation function, which is expanded on in the next section.
@@
The output is the $m$’th element of the state of the reservoir at the next time step, $k+1$. It is switched into the appropriate output branch by a $1:N$ switch and the output is given an appropriate delay and split before both being sent to a detector and back into the reservoir for the next time step. This design utilizes only a single nonlinear activation element, allowing for high scalability. After the necessary delays to line up all of the output vectors, and at the end of all of the sub-steps, the state vector has been successfully updated and fed back into the reservoir for the next time step. After collecting all of the state vectors in the detection steps, we use post-training according to Equation (2) to achieve the required estimated output in response to the time-series inputs $u_{k}$.
@@
This design allows for the choice of arbitrary matrix elements for the matrix $A$ and vector $B$, including negative elements in dBm units through the use of the appropriate attenuation before each Brillouin amplifier. If chosen according to the convergence properties of an ESN, this design can be used to implement an arbitrary ESN, which is a universal family of reservoir computers [17]. This means that given a desired accuracy for a task (such as prediction), this method can be used to realize an optical hardware that achieves that accuracy. The proposed design also allows for scalability, as the number of nodes is increased simply by increasing the number of parallel fibers running in the optical fiber reservoir. Although, scalability is naturally limited by available resources and energy.
An important consideration is the energy consumption of the system. While an arbitrary ESN can be implemented with this architecture, the power flowing through the system can approach unfeasible amounts. In order to limit this, the scaling criteria on the amplifiers and nonlinear activation is made stringent to ensure that the maximum value of $x_{k}$ in the system is $31$ dBm. The maximum pump power for the nonlinear activation required in this scenario is $3.2$ W for a pump splitting ratio of 1000. The pump splitting ratio is discussed further in Section IV. While this results in a slight degradation in performance, we must accept this tradeoff to keep the power budget low.
[…]
The delay line solution adds a different kind of hardware complexity despite reducing the required number of Brillouin amplifiers. While this is a drawback of using a fiber-based approach, we consider this fiber-based architecture to be a preliminary step towards an on-chip all optical architecture. Choudhary et al. [28], and Yu et al. [29] have shown >50>50 dB of Brillouin amplifier gain in photonic chip and >8>8 dB of gain in thin-film lithium niobate solutions respectively, showing promise that a high-gain and tunable Brillouin amplifier for the construction of a similar on-chip optical ESN may be feasible in the near future. In addition, the on-chip photonic-circuit implementation will allow multiple nonlinear activating Brillouin amplifiers in parallel, obviating the need for delay lines. This future photonic-circuit implementation will ensure the scalability of the machine size.
On-chip implementations will also simplify the phase locking requirement. These systems have much lower phase fluctuations than fiber systems due to the solid on-chip structure, and approaches to combine coherent beams in Silicon On Insulator (SOI) fabrications, such as the method demonstrated by Jiao et al. [30], show promise for this ESN technology to be brought to an on-chip architecture.
IV Results and Discussion
@@
(a) Software (Sigmoid) Sine Square Classification, NMSE=0.14.
(b) Optical Sine Square Classification, NMSE=0.12.
Good performance in machine learning hinges on the ability to approximate nonlinear input–output relationships and capture complex temporal dynamics. According to the Stone–Weierstrass theorem [31], the polynomial algebra property is central to achieving universality. Therefore, we selected three distinct tasks to evaluate these critical capabilities. They are the sine-square classification test, prediction of the NARMA10 system, and the Mackey-Glass time series prediction. Each task evaluates mainly the reservoir computer nonlinearity, polynomial algebra capabilities, and dynamic memory retention, respectively. Additionally, we address the nonlinear properties of the SBS interaction and, thus, the optimization of the activation function. To achieve good results in the tasks mentioned above, we require strong enough nonlinearity, while maintaining the convergence conditions of the ESN. We vary the value of the $gL$ product as well as the pump splitting ratio with these constraints in mind to optimize the performance of the ESN in the selected benchmark tests.
@@
For each test, we analyze the effect of the nonlinear strength of the SBS interaction, given by the $gL$ product value of the medium and the pump splitting ratio for the self-pumped nonlinear activation element, which refers to the ratio of the pump power to the input power. Here, $g$ represents the Brillouin gain and $L$ represents the interaction length. These values appear in Equation (7). The $gL$ product decides the nonlinear shape and the parking point of the activation function, as shown in Figure 1. The pump splitting ratio independently changes the parking point of the activation function, which allows us to optimize it for greater performance. It is important to note however that increasing the splitting ratio increases the power budget of the system. The parking point of the tunable activation function moves between -5 dBm to -10 dBm for the range of gL values that we have used for testing in this work, as shown in Figure 1. This means that, around the parking points, the pump power will be 100300100-300 mW for a splitting ratio of 1000. Since we have allowed for a maximum output $x_{k}$ value of $31$ dBm, the maximum pump power required in this scenario is $3.2$ W, which is quite a feasible power budget for the nonlinear activation elements. Better performance can generally be achieved for larger splitting ratios, which means that the best performance is restricted by the energy availability.
(a) Average NMSE vs. Pump Splitting ratio for various gLgL, $N=10$.
(b) Average NMSE vs. Pump Splitting ratio for various NN, $gL=12$.
(c) Average NMSE vs. gLgL values for various NN, $\mathrm{Pump/Input}=1000$.
We also evaluate the performance as the number of nodes $N$ is increased, which shows the effect of scaling the ESN. These behaviors are task-dependent and the exact values of these parameters are subject to change with randomization of the ESN and inputs but, by extracting a range of top-performing parameters across the board, we attempt to design the most generally applicable SBS activation function and optical ESN. We do this by analyzing the behavior for 250 trials of different $A$ and $B$ matrices. These investigations are expanded on below.
In order to simulate the proposed architecture for the matrix-vector multiplication, the data at the input and output is always represented in dBm units, which allows access to negative elements as discussed in the previous section. However, the input dBm data is always converted to W units before the multiplication step. This must be done in order to accurately simulate the behavior of optical components such as VOAs and Brillouin amplifiers. The output is converted back to dBm for the detection phase, ensuring that the data is always represented in dBm units.
IV.1 Sine Square Classification
@@
We evaluate the effect of varying activation function parameters on the NMSE for this test. We have computed the average NMSE over 250 variations of the $A$ and $B$ matrices to see clear trends in the NMSE. From Figures 4 (a) and (b), NMSE decreases consistently with increasing pump splitting ratios, indicating improved nonlinear activation efficiency. Figure 4 (c) shows that for a fixed pump splitting ratio as well, the performance nearly saturates at lower $N$ values for $gL$ products in the range of $3-7\;W^{-1}$, with slight (<0.2<0.2) improvements in NMSE for higher $gL$ values in the case of larger $N$ values. We see that for $N=4$ nodes, the trends are difficult to see and the NMSE values are much higher. This is because the fitting is difficult for a smaller number of nodes. Therefore, we find for this task that $N=5$ and above are required for reasonable performance, with an increase in $N$ taking the performance closer and closer to saturation near $\mathrm{NMSE}\approx 0$. We also see, from Figure 4 (a) that there is a different optimum point in the performance for each $gL$ value. The most probable reasoning for this behavior is that both the pump splitting ratio and the $gL$ value are together responsible for the optimal parking point of the activation function, resulting in local minima.
@@
The values reported by Slinkov et. al [22] allow the possibility of $gL$ values from 373-7 with fiber lengths of 363-6m at reasonable threshold powers of $1.26$ W which are very practical lengths of fiber. Experimental demonstrations have shown much higher levels of nonlinearity in the right materials [33, 28]. This shows promise for the translation of this architecture into an on-chip implementation, which has the benefit of improved scalability while maintaining the other advantages of this fiber-based approach.
IV.2 NARMA10
(a) Average NMSE vs. Pump Splitting ratio for various gLgL, $N=10$.
(b) Average NMSE vs. Pump Splitting ratio for various NN, $gL=12$.
(c) Average NMSE vs. gLgL values for various NN, $\mathrm{Pump/Input}=1000$.